# SPECIAL RELATIVITY

### v9

@@ 3.4 Motion of a Light Source With Respect to the Inertial Frame
We now consider the case when the light source in the lab frame is accelerated from rest up to velocity $v$ along the $x$-axis. In other words, we consider so-called active (physical) boost. When we state that the source undergoes an acceleration, and the inertial observer (with measuring devices) does not, the acceleration means acceleration relative to the fixed stars. Any acceleration relative to the fixed stars (i.e. any active boost of velocity) has an absolute meaning.
In the case of an active boost we are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system. A fundamental question to ask is whether our lab clock synchronization method depends on the state of motion of the light source or not. The answer simply fixes a convention. The simplest method of synchronization consists in keeping, without changes, the same set of uniformly synchronized clocks used in the case when the light source was at rest, i.e. we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest). This choice is usually the most convenient one from the viewpoint of connection to laboratory reality. This synchronization convention preserves simultaneity and is actually based on the absolute time (or absolute simultaneity) convention.
Absolute simultaneity can be introduced in special relativity without affecting neither the logical structure, no the (convention-independent) predictions of the theory. We begin with the metric as true measure of space-time intervals for an inertial observer with coordinates $(t,x,y,z)$. Using the Galilean boost: we substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain
@@
This metric describes the electrodynamics of the moving light source with viewpoint of the inertial observer measurements. From the formula itself it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x$ and time $t$ of the measured device at rest in the inertial frame is the same as the measured result of this configuration at point $x-vt$ and time $t=0$.
@@
To understand the physical interpretation of the Eq. 11, it is important to note that there are infinitely many ways to label points in space-time by coordinates. The different coordinate systems are connected by coordinate transformations. It is always possible to create a new frame of reference by relabeling coordinates, and then discussing physical phenomena in terms of the new coordinate labels - a passive (coordinate) transformation. For example, it is always possible to create a so-called comoving coordinate system in the lab frame, and then discussing radiation from the moving source in therm of the new (comoving) coordinate labels. In the comoving coordinate system, fields are expressed as a function of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$. The variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ by means of a passive Galilean transformation, so that fields can be written in terms of $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$. After the passive transformation, the Cartesian coordinates of the source transform as $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z$. This transformation completes with the invariance of simultaneity, $t^{\prime}=t$. The transformation of time and spatial coordinates of any event has the form of a Galilean transformation. One must take into account that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
@@
The equivalence of the active and passive pictures is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount. Obviously, there is no a priory guarantee that the world behave like this. For instance, Newton mechanics is not invariant under passive rotation as demonstrated Newton’s bucket experiment. The empirical principle implies that within a single inertial frame every passive Galilean boost is postulated to corresponds to an active Galilean boost. According to this equivalence of passive and active boosts, in our case of interest the Maxwell’s equations always valid in the comoving frame. In the comoving frame, fields are expressed as a function of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$. The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. However, the variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ by means of a Galilean transformation, so that fields can be written in terms of $x,y,z,t$. From the Galilean transformation $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z,~{}t^{\prime}=t$, after partial differentiation, one obtains $\partial/{\partial t}=\partial/{\partial t^{\prime}}-v\partial/{\partial x^{\prime}}$, $\partial/{\partial x}=\partial/{\partial x^{\prime}}$. Hence the wave equation transforms into
@@
The solution of this equation $F[x-(c+v)t]+G[x+(-c+v)t]$ is the sum of two arbitrary functions, one of argument $x-(c+v)t$ and the other of argument $x+(-c+v)t$. Here we obtained the solution for waves which move in the $x$ direction by supposing that the field does not depend on $y$ and $z$. The first term represents a wave traveling forward in the positive $x$ direction, and the second term a wave traveling backwards in the negative direction.
We conclude that the speed of light emitted by a moving source measured in the lab frame $(t,x)$ depends on the relative velocity of source and observer, in our example $v$. In other words, the speed of light is compatible with the Galilean law of addition of velocities. In fact, the coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c+v$ in the positive direction, and $dx/dt=-c+v$ in the negative direction. The reason why it is different from the electrodynamics constant $c$ is due to the fact that the clocks are synchronized following the absolute time convention. One fixed, such convention should be applied and kept in a consistent way for both dynamics and electrodynamics. In fact, the light velocity transforms as a particle velocity in the absolute time coordinatization as it mast be.
@@ 3.5 Discussion
So far we have considered the Galilean transformation of the electrodynamics equations. We would like to emphasize a very important difference between fundamental and phenomenological theory. For example, we can characterize Newton’s equation in the co-moving frame as a phenomenological law. The microscopic interpretation of the inertial mass of a particle is not given and the rest mass is introduced in an ad hoc manner. It can be said that the particle has some inner variables - that we do not yet know about. Any phenomenological law, which is valid in the Lorentz rest frame, can be embedded in the four dimensional space-time only by using Lorentz coordinatization. Now we wish to consider the electrodynamics. It is important to stress at this point that there is no machinery (i.e. inner variables) behind the laws of electrodynamics. It is obvious that electromagnetic fields are fundamental and the electrodynamics theory meets all requirements of the theory of relativity. Consequently, one can use any synchronization to describe the electromagnetic field dynamics. The variety of methods of description physical systems makes it possible to select for each problem the representation most suitable for its solution. It is always possible to create a new frame of reference by relabeling coordinates, and then discussing physical phenomena in terms of the new coordinate labels - a passive transformation. An active transformation of a physical system is its motion, i.e., a variation in its characteristics under the effect of some internal or external interactions. We are dealing with motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
@@
Now let us discuss correspondence between passive and active transformations. A comparison between the special relativity and the Newton theory might help here. The formalism of non-relativistic Newton’s theory is associated with absolute space structure. However, there is relativity of the uniform motions and we cannot determine the particle absolute velocity. Passive Galilean boosts are dynamical symmetries of Newtonian mechanics. The equivalence of the active and passive Galilean boost is due to the fact that moving system one way is equivalent to moving the coordinate system the other way by an equal amount as described by the famous (Galileo) ship experiment . The reason of this (Galilean group) symmetry is simply that equations of motion do not explicitly depend on velocity. The internal dynamics of the system does not change under the Galilean boost. In the same time, Newtonian mechanics is not invariant under passive rotation. In fact, the form of equations of motion change due to pseudo-force terms (4). The Newton bucket argument by Newton was based on comparison of two different situations: a vessel of water at rest, vs the vessel after it has gradually been rotated from rest to a state of uniform angular velocity. The argument against the equivalence of the active and passive rotation is based on notion that in both cases there is no relative motion between the water and the bucket, while the shape of the water turns out to be different. The passive rotation transformation is no more than a change of representation. There is no relativity of accelerations and we can determine the particle absolute acceleration in the inertial frame. The absolute acceleration means acceleration relative to the fixed stars.
@@
Let us now move to consider correspondence between active and passive transformations in special relativity. Within in single inertial frame the kinematical symmetry is corresponds to dynamical symmetry and the symmetry principle (i.e. equivalence of the active and passive Galilean boost) is satisfied. It should be note that passive Galilean boost within a single inertial frame is simply another parametrization of the observations of the inertial observer. According to passive transformation ($x\to x-vt$, $t\to t$), the motion of fixed stars with respect to the observer and his devices not changes. It is important to note that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
@@
In the Chapter 5 we will consider the problem of aberration of light in accelerated systems on the basis of the theory of special relativity. This problem can be adequately treated only by an approach which uses the absolute time coordinatization. In order to write electrodynamics laws in an accelerated frame, we define the metric of the accelerated frame using the inverse Galilean boost. We want to emphasize that in the case of an inverse Galilean boost of velocity we consider the effect of interaction which defined in terms of acceleration motion relative to the fixed stars. Any acceleration relative to the fixed stars has an absolute meaning.
@@
This means that the passive Galilean transformation within a single inertial frame is quite distinct from the (passive) inverse Galilean transformation, $x\to x_{n}+vt$, $t\to t_{n}$, i.e acceleration with respect to fixed stars of an observer (with his measuring instruments) from inertial frame to the accelerate frame. One must take into account that $(t_{n},x_{n},y_{n},z_{n})$ are coordinates of the rule clock structure at rest in the accelerated frame. Such transition of the coordinates of the light source does not stand for a change of representation, but for a change in the absolute motion of the light source with respect to the fixed stars. The fact that in the real process of transmission to an accelerated (with respect to the fixed stars) frame the observer will experience the inertial force, is not accounted in the standard treatments. We argue that reciprocity of the inertial frames, which is considered in textbooks as relativistically correct, is at odds with special relativity. The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and accelerated reference systems are not equitable.
@@
The formalism of relativistic physics is associated with absolute space-time structure. There is no relativity of the uniform motions and we can determine the system absolute velocity in the initial inertial frame. The reason of this asymmetry is simply that electrodynamics equations are fundamental and explicitly depend on velocity. According to the special relativity, there is a real difference between accelerated (with respect to the fixed stars) inertial frame and the inertial frame without accelerational history and this difference is closely associated with the inertial forces. It should be note that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame. Nevertheless, the integral effect of inertial forces completely determined the problem. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer. We can treat a uniformly accelerating frame as if it were an in inertial frame with the addition of a uniform pseudo-gravitational field. This is basic content of equivalence principle. The principle of equivalence can be applied to solve non-inertial kinematics problems with dynamics methods. For instance, when the accelerated system starts moving with constance velocity the pseudo-gravitational field is zero. For the gravitational potential difference between the two frames we get (in the simplified second order approximation) $\phi=-v^{2}/2$. The fact that the inertial observer always turns out to have aged more that accelerated one, in the equivalence principle analysis is elementary consequence of the slowly down of the rate of the processes in the pseudo-gravitational potential. In fact, for clock at rest in the accelerated frame we have $dt_{prop}=dt_{n}(1+\phi/c^{2})=dt_{n}[1-v^{2}/(2c^{2})]$. In other words, the time dilation effect is absolute, but not relative.
@@
There is another way to solve the problem. From a space-time geometrical viewpoint, the explanation consists in using the fact that metric tensor must be continuous quantity. When the accelerated system starts moving with constant velocity the procedure of smooth tailoring of metric tensor can be performed. The (non-diagonal) Langevin metric is found by matching the accelerated frame and initial inertial frame metric tensors. Many experts who learned the theory of relativity using textbooks will find this statement disturbing. Indeed, according to textbooks, the problem is complete symmetric and both frames are equivalent. The popular belief remains that coordinates of these frames should be related by the Lorentz transformations and the metric in both frames will be diagonal, although it is an exercise in elementary algebra to prove discontinuity of metric tensor. With mathematical viewpoint, we must conclude that when authors of textbooks are dealing with transformation and this transformation does not preserve continuity of the metric tensor there are no symmetry between inertial frames.
The initial inertial frame is not physically identical to the Newton’s absolute space. The special relativity taught us to think in terms of pseudo-Euclidean space-time geometry. In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time. We will discuss this subject further in the Chapter 5, Chapter 9.

### v10

@@ 3.3 Motion of a Light Source With Respect to the Inertial Frame
We now consider the case when the light source in the lab frame is accelerated from rest up to velocity $v$ along the $x$-axis. In other words, we consider a so-called active (physical) boost. When we state that the source undergoes an acceleration, and the inertial observer (with measuring devices) does not, the acceleration means acceleration relative to the fixed stars. Any acceleration relative to the fixed stars (i.e. any active boost of velocity) has an absolute meaning.
In the case of an active boost, we are dealing with the motion of the same physical system, evolving in time and treated from the point of view of the same reference system. A fundamental question to ask is whether our lab clock synchronization method depends on the state of motion of the light source or not. The answer simply fixes a convention. The simplest method of synchronization consists of keeping, without changes, the same set of uniformly synchronized clocks used in the case when the light source was at rest, i.e. we still enforce the clock transport synchronization ( or Einstein synchronization which is defined in terms of light signals emitted by the dipole source at rest). This choice is the most convenient one from the viewpoint of connection to laboratory reality. This synchronization convention preserves simultaneity and is based on the absolute time (or absolute simultaneity) convention.
Absolute simultaneity can be introduced in special relativity without affecting neither the logical structure nor the (convention-independent) predictions of the theory. We begin with the metric as the true measure of space-time intervals for an inertial observer with coordinates $(t,x,y,z)$. Using the Galilean boost: we substitute $x\to x-vt$, while leaving time unchanged into the Minkowski metric $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$ to obtain
@@
This metric describes the electrodynamics of the moving light source from the viewpoint of the inertial observer measurements. From the formula itself, it is easy to see that the result of measurement of the electromagnetic field configuration in the moving system in terms of coordinate $x$ and time $t$ of the measured device at rest in the inertial frame is the same as the measured result of this configuration at point $x-vt$ and time $t=0$.
@@
To understand the physical interpretation of the Eq. 11, it is important to note that there are infinitely many ways to label points in space-time by coordinates. The different coordinate systems are connected by coordinate transformations. It is always possible to create a new frame of reference by relabeling coordinates, and then discussing physical phenomena in terms of the new coordinate labels - a passive (coordinate) transformation. For example, it is always possible to create a so-called comoving coordinate system in the lab frame, and then discuss radiation from the moving source in therm of the new (comoving) coordinate labels. In the comoving coordinate system, fields are expressed as a function of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$. The variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ using a passive Galilean transformation so that fields can be written in terms of $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$. After the passive transformation, the Cartesian coordinates of the source transform as $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z$. This transformation completes with the invariance of simultaneity, $t^{\prime}=t$. The transformation of time and spatial coordinates of any event has the form of a Galilean transformation. One must take into account that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
@@
The equivalence of the active and passive pictures is because moving system one way is equivalent to moving the coordinate system the other way by an equal amount. Obviously, there is no a priory guarantee that the world behave like this. For instance, Newton’s mechanics is not invariant under passive rotation as demonstrated Newton’s bucket experiment. The empirical principle implies that within a single inertial frame, every passive Galilean boost is postulated to correspond to an active Galilean boost. According to this equivalence of passive and active boosts, in our case of interest, Maxwell’s equations are always valid in the comoving frame. In the comoving frame, fields are expressed as a function of the independent variables $x^{\prime},y^{\prime},z^{\prime}$, and $t^{\prime}$. The electric field $\vec{E}^{\prime}$ of an electromagnetic wave satisfies the equation $\Box^{\prime 2}\vec{E}^{\prime}=\nabla^{\prime 2}\vec{E}^{\prime}-\partial^{2}\vec{E}^{\prime}/\partial(ct^{\prime})^{2}=0$. However, the variables $x^{\prime},y^{\prime},z^{\prime},t^{\prime}$ can be expressed in terms of the independent variables $x,y,z,t$ using a Galilean transformation, so that fields can be written in terms of $x,y,z,t$. From the Galilean transformation $x^{\prime}=x-vt,~{}y^{\prime}=y,~{}z^{\prime}=z,~{}t^{\prime}=t$, after partial differentiation, one obtains $\partial/{\partial t}=\partial/{\partial t^{\prime}}-v\partial/{\partial x^{\prime}}$, $\partial/{\partial x}=\partial/{\partial x^{\prime}}$. Hence the wave equation transforms into
@@
The solution of this equation $F[x-(c+v)t]+G[x+(-c+v)t]$ is the sum of two arbitrary functions, one of argument $x-(c+v)t$ and the other of argument $x+(-c+v)t$. Here we obtained the solution for waves that move in the $x$ direction by supposing that the field does not depend on $y$ and $z$. The first term represents a wave traveling forward in the positive $x$ direction, and the second term represents a wave traveling backward in the negative direction.
We conclude that the speed of light emitted by a moving source measured in the lab frame $(t,x,y,z)$ depends on the relative velocity of the source and observer, in our example $v$. In other words, the speed of light is compatible with the Galilean law of addition of velocities. In fact, the coordinate velocity of light parallel to the $x$-axis is given by $dx/dt=c+v$ in the positive direction, and $dx/dt=-c+v$ in the negative direction. The reason why it is different from the electrodynamics constant $c$ is due to the fact that the clocks are synchronized following the absolute time convention. One fixed, such convention should be applied and kept in a consistent way for both dynamics and electrodynamics. In fact, the light velocity transforms as a particle velocity in the absolute time coordinatization as it must be.
@@ 3.4 Discussion
So far we have considered the Galilean transformation of the electrodynamics equations. We would like to emphasize a very important difference between fundamental and phenomenological theory. For example, we can characterize Newton’s equation in the co-moving frame as a phenomenological law. The microscopic interpretation of the inertial mass of a particle is not given and the rest mass is introduced in an ad hoc manner. It can be said that the particle has some inner variables - that we do not yet know about. Any phenomenological law, that is valid in the Lorentz rest frame, can be embedded in the four-dimensional space-time only by using Lorentz coordinatization. Now we wish to consider the electrodynamics. It is important to stress at this point that there is no machinery (i.e. inner variables) behind the laws of electrodynamics. Electromagnetic fields are fundamental and the electrodynamics theory meets all requirements of the theory of relativity. Consequently, one can use any synchronization to describe the electromagnetic field dynamics. The variety of methods of description of physical systems makes it possible to select for each problem the representation most suitable for its solution. It is always possible to create a new frame of reference by relabeling coordinates, and then discussing physical phenomena in terms of the new coordinate labels - a passive transformation. An active transformation of a physical system is its motion, i.e., a variation in its characteristics under the effect of some internal or external interactions. We are dealing with the motion of the same physical system, evolving in time and treated from the point of view of the same reference system.
@@
Now let us discuss the correspondence between passive and active transformations. A comparison between special relativity and Newton’s theory might help here. The formalism of non-relativistic Newton’s theory is associated with absolute space structure. However, there is relativity of the uniform motions and we cannot determine the particle’s absolute velocity. Passive Galilean boosts are dynamical symmetries of Newtonian mechanics. The equivalence of the active and passive Galilean boost is because that moving system one way is equivalent to moving the coordinate system the other way by an equal amount as described by the famous (Galileo) ship experiment. The reason for this (Galilean group) symmetry is simply that equations of motion do not explicitly depend on velocity. The internal dynamics of the system does not change under the Galilean boost. At the same time, Newtonian mechanics is not invariant under passive rotation. The form of equations of motion change due to pseudo-force terms  [^13] For a general discussion of correspondence between passive and active transformations we suggest reading the paper [21].. The Newton bucket argument by Newton was based on a comparison of two different situations: a vessel of water at rest, vs the vessel after it has gradually been rotated from rest to a state of uniform angular velocity. The argument against the equivalence of the active and passive rotation is based on the notion that in both cases there is no relative motion between the water and the bucket, while the shape of the water turns out to be different. The passive rotation transformation is no more than a change of representation. There is no relativity of accelerations and we can determine the particle’s absolute acceleration in the inertial frame. The absolute acceleration means acceleration relative to the fixed stars.
@@
Let us now move to consider the correspondence between active and passive transformations in special relativity. Within a single inertial frame, the kinematical symmetry corresponds to dynamical symmetry and the symmetry principle (i.e. equivalence of the active and passive Galilean boost) is satisfied. It should be noted that passive Galilean boost within a single inertial frame is simply another parametrization of the observations of the inertial observer. According to passive transformation ($x\to x-vt$, $t\to t$), the motion of fixed stars with respect to the observer and his devices does not change. It is important to note that $(t,x,y,z)$ are coordinates of the rule clock structure at rest in the inertial frame.
@@
In Section 5 we will consider the problem of aberration of light in accelerated systems based on the theory of special relativity. This problem can be adequately treated only by an approach that uses absolute time coordinatization. To write electrodynamics laws in an accelerated frame, we define the metric of the accelerated frame using the inverse Galilean boost. We want to emphasize that in the case of an inverse Galilean boost of velocity, we consider the effect of interaction which is defined in terms of acceleration motion relative to the fixed stars.
@@
This means that the passive Galilean transformation within a single inertial frame is quite distinct from the (passive) inverse Galilean transformation, $x\to x_{n}+vt$, $t\to t_{n}$, i.e acceleration with respect to fixed stars of an observer (with his measuring instruments). One must take into account that $(t_{n},x_{n},y_{n},z_{n})$ are coordinates of the rule clock structure at rest in the accelerated frame. Such transition of the coordinates of the light source does not stand for a change of representation but for a change in the absolute motion of the light source (and rule-clock structure) with respect to the fixed stars. The fact that in the real process of transmission to an accelerated (with respect to the fixed stars) frame the observer will experience the inertial force, is not accounted for in the standard treatments. We argue that the reciprocity of the inertial frames, which is considered in textbooks as relativistically correct, is at odds with special relativity. The arguments concerning the relativity of motion in our case of interest cannot be applied, since the inertial and accelerated reference systems are not equitable.
@@
According to special relativity, there is a real difference between the accelerated (with respect to the fixed stars) inertial frame and the inertial frame without accelerational history, and this difference is closely associated with the inertial forces. It should be noted that the duration of the acceleration period has a negligible effect on the anisotropy in the accelerated frame. Nevertheless, the integral effect of inertial forces completely determined the problem. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer. We can treat a uniformly accelerating frame as if it were an inertial frame with the addition of a uniform pseudo-gravitational field. This is the basic content of the equivalence principle. The principle of equivalence can be applied to solve non-inertial kinematics problems with dynamics methods.
@@
There is another way to solve the problem. The formalism of relativistic physics is associated with absolute space-time structure. There is no relativity of the uniform motions and we can determine the system’s absolute velocity in the initial inertial frame. From a space-time geometrical viewpoint, the explanation consists in using the fact that the metric tensor must be a continuous quantity. When the accelerated system starts moving with constant velocity the procedure of smooth tailoring of the metric tensor can be performed. The (non-diagonal) Langevin metric is found by matching the accelerated frame and initial inertial frame metric tensors. Many experts who learned the theory of relativity using textbooks will find this statement disturbing. Indeed, according to textbooks, the problem is completely symmetric and both frames are equivalent. The popular belief remains that the coordinates of these frames should be related by the Lorentz transformations and the metric in both frames will be diagonal. However, after diagonalization, the metric tensor in the accelerated frame must abruptly change from the Langevin value to the Minkowski value. Therefore, contrary to the textbook view, Lorentz transformation does not preserve the continuity of the metric tensor. We will discuss this subject further in Chapter 5.